OpenStudy (anonymous):

What is the sum of the measures of the interior angles of a regular polygon where a single exterior angle measures 72°?

OpenStudy (anonymous):

360/72 = 5 (sides)

OpenStudy (cwrw238):

one internal angle = 180-72 = 108 degrees and there are 5 of these